OpenStudy (anonymous):

x-3y=-6 -x+3y=6 a) the system has no solution b) the system has a unique solution (x,y)=(_,_) c) the system has infinitely many solutions. they must satisfy the following equation y=

OpenStudy (f_jayyy):

If you have both X and Y in an equation there is no solution, so the answer would be A

OpenStudy (anonymous):

thank you @F_Jayyy

OpenStudy (texaschic101):

I disagree....if at the end, both sides are equal, then there are infinite solutions

OpenStudy (texaschic101):

when adding these equations, you end up with 0 = 0....equal.

OpenStudy (anonymous):

ok now im confused.. who is right ? can either of you explain how you got the answer

OpenStudy (f_jayyy):

But it is not 0 because 3 would go into 6 at least twice

OpenStudy (texaschic101):

I am trying to think of a way to explain this.... x - 3y = -6 -x + 3y = 6 ----------add 0 + 0 = 0 this implies infinite solutions if it would have ended like : 0 = 3...or something unequal, then there would be no solutions if it would have ended like : x = 3...or y = 2...then there would be 1 solution